# A Two-Stage Stochastic Model for Road-Rail Intermodal Freight Transportation Under Demand and Capacity Uncertainty

### v1

@@ 5.3.3 Impact of Train Capacity Constraints on Demand and Total Cost
Train capacity constraints play a crucial role in intermodal transport, especially under stochastic demand and varying train capacity. To maintain computational efficiency, this experiment adopts a streamlined setup. The study examines a network of 4 freight trains operating under 8 demand scenarios across 22 intermodal hubs, with a total demand of 1013 containers, averaging 32 per train. This setup maintains the key stochastic elements while focusing on the analysis of the trade-offs between capacity expansion, cost reduction, and service reliability. The results, summarized in Table 3, shows how the capacity of the train influences unmet demand and costs, directly addressing the sixth policy question by identifying the threshold beyond which additional capacity offers diminishing cost savings.
@@
Effect of Train Capacity on System Performance
Train capacity constraints directly impact cost and demand satisfaction, with increasing capacity leading to significant cost reductions and improved service levels. However, beyond a certain threshold, additional capacity yields diminishing returns. As shown in Table 3, increasing capacity from 4 to 8 containers per station significantly lowers costs and unmet demand, but further expansion to 10 containers results in only marginal improvements.
This trend is driven by the interaction between the train capacity constraint and the model’s objective:
  • Capacity Bottlenecks at Lower Levels: When capacity is low, the constraint (7) restricts container allocations, leading to unmet demand penalties. Increasing capacity alleviates these bottlenecks, improving system efficiency.
  • Scenario-Specific Demand Surges: With 8 demand scenarios, fluctuating demand levels create imbalances. At lower capacities, trains struggle to absorb peak demands, causing demand shortages. As capacity expands, the model gains more flexibility, reducing cost volatility.
  • Trade-offs in Cost Optimization: The model balances minimizing unmet demand penalties and optimizing transportation efficiency. While increasing capacity initially reduces penalties, beyond 8 containers per station, cost savings plateau as the system reaches near-optimal efficiency.
Interplay Between Cost and Unmet Demand
The relationship between total cost and unmet demand is further illustrated in Figure 6, which provides a visual representation of how capacity constraints shape system performance. At low capacity levels, the high unmet demand significantly contributes to the total system cost due to penalty terms associated with demand shortfalls. As capacity increases, unmet demand declines rapidly, leading to substantial cost reductions. However, beyond 8 containers per station, additional capacity improvements yield only marginal cost savings. This behavior suggests that while increasing capacity is effective in reducing costs initially, other factors such as supply constraints, competition for train slots, and network-wide operational limits begin to dominate decision-making at higher capacity levels.
The standard deviation of unmet demand across trains provides additional insights into how evenly shortages are distributed. At low capacity (4 containers per station), the system experiences significant demand imbalances, with the maximum unmet demand reaching 18 containers for a single train, while the minimum unmet demand remains 10 containers. This suggests that certain routes or stations experience significantly more congestion than others.
@@
As capacity increases, unmet demand becomes more evenly distributed, with both its maximum value and standard deviation decreasing. By capacity 10, the maximum unmet demand is just 2 containers, and the standard deviation drops to 0.67, reflecting a more balanced use of available transport resources. This stabilization suggests that beyond a certain threshold, train capacity is no longer the primary bottleneck, other operational constraints, such as schedule synchronization, become more dominant in affecting system’s performance.
@@
The following key takeaways emerge from the results:
  • Increasing train capacity significantly improves demand satisfaction, but beyond a certain threshold (8 containers per station in our case), further cost reductions are minimal due to other constraints becoming binding (e.g., supply availability and schedule alignment).
  • At lower capacities, demand imbalances are more pronounced, leading to high unmet demand variability across trains. This variability diminishes at higher capacities, indicating a more balanced network.
  • The model balances the trade-off between cost and demand fulfillment, minimizing penalties from unmet demand while ensuring train utilization remains efficient.
  • Network performance is highly sensitive to the available capacity per station, which is often influenced by external factors such as demand fluctuations and competition from other companies utilizing the same train services. Given these uncertainties, planning should focus on adapting to dynamic availability.
These insights reinforce the importance of strategic capacity planning in intermodal transport. As spot capacity availability increases, the system gains more flexibility in allocating containers, reducing the risk of unmet demand. However, relying solely on capacity expansion may not be the most cost-effective approach. Instead, decisions should prioritize adaptive scheduling, dynamic capacity allocation, and improved coordination among stakeholders to enhance system resilience and maximize efficiency under uncertain capacity conditions.
@@ 5.3.4 Sensitivity of Cost Metrics to Risk Preferences and Scenario Variability
A key challenge in stochastic intermodal transportation planning is ensuring decision robustness across cost-risk trade-offs and uncertainty scenarios. While prior analyses have examined the impact of $\lambda$ and $\alpha$ on total cost and CVaR, it is equally important to assess how these parameters influence cost variability and model stability. This section investigates, the divergence between objective value (OBJ) and expected total cost ($E(TC)$), cost variability across scenarios, and the stability of cost structures under different risk settings. Figures 7, 9, and 8 provide a visualization of these relationships.
@@
This analysis addresses the first policy question by illustrating how $\lambda$ and $\alpha$ affect cost predictability, resilience, and financial trade-offs. It also provides insights into the seventh policy question, highlighting at what point additional risk mitigation yields diminishing returns. Additionally, the findings contribute to the third policy question by demonstrating how different risk aversion levels impact cost robustness under varying demand conditions.
Discrepancy Between Objective Value and Expected Total Cost
A key observation from Figure 7 is the increasing gap between OBJ and $E(TC)$ as $\lambda$ rises, particularly at higher $\alpha$ values. At low $\lambda$, the two metrics remain closely aligned, indicating that when risk aversion is minimal, the model prioritizes cost efficiency with little regard for extreme scenarios. However, as $\lambda$ increases, OBJ becomes increasingly dominated by CVaR penalties, reflecting the fundamental trade-off between cost minimization and resilience to worst-case outcomes.
Scenario Variability and Cost Stability
Figure 8 shows that at low $\lambda$ values, total cost fluctuates significantly across scenarios, as the model optimizes for average cost rather than stability. As $\lambda$ increases, variability decreases, leading to more predictable cost structures. However, beyond $\lambda\approx 0.8$, further risk aversion yields diminishing reductions in cost fluctuations, suggesting that excessive risk mitigation comes at a steep cost with limited additional benefits.
Insights from CVaR Sensitivity Analysis
Figures 9 highlight how risk preferences influence extreme cost outcomes. While increasing $\lambda$ substantially reduces CVaR, the effect is more pronounced at higher confidence levels ($\alpha=0.90$) compared to lower ones. Notably, the gains from increasing $\lambda$ taper off beyond $\lambda\approx 0.5$, reinforcing that moderate risk aversion ($\lambda$ between 0.4-0.6) strikes an effective balance between mitigating extreme costs and maintaining overall cost efficiency.
@@
Decision Robustness and Practical Implications
From a practical decision-making standpoint, these results reveal crucial insights into how risk preferences influence operational consistency and cost predictability:
  • For cost-minimization strategies ($\lambda\leq 0.3$): High cost variability across scenarios, more aggressive cost allocation, which may result in unanticipated financial risks when extreme-cost scenarios materialize, and suitable for stakeholders prioritizing efficiency over stability, particularly in low-volatility environments.
  • For balanced risk strategies ($\lambda\approx 0.5$): Moderate cost stability with significant CVaR reduction, lower variability in total costs, providing a predictable financial structure without excessive cost premiums, and suitable for stakeholders seeking a risk-aware but cost-efficient approach.
  • For highly risk-averse strategies ($\lambda\geq 0.8$): minimal cost fluctuations, leading to greater financial predictability, substantial increase in total cost, reducing cost efficiency in favor of extreme-risk protection, and best suited for industries with high exposure to financial uncertainty, such as global supply chains with high operational risk.
Figure 9: Visualization of CVaR values under varying $\lambda$ and $\alpha$ highlighting the sensitivity of CVaR to changes in $\lambda$ and $\alpha$.
Broader Implications and Future Considerations
These insights directly address the policy questions on cost-risk balancing and decision adaptability, demonstrating that the optimal choice of $\lambda$ and $\alpha$ depends on the decision-making priorities of different stakeholders. For future research, exploring interactions between train capacity constraints, emissions penalties, and risk-aversion strategies could further enhance the practical applicability of these findings. Additionally, expanding the scenario space to include real-world disruptions, such as terminal congestion and supply chain bottlenecks, would provide a more comprehensive understanding of risk-sensitive transportation planning.
@@ 5.3.5 Analyzing the Influence of Emissions Cap on Model Performance
The incorporation of emissions constraints in optimization models is a crucial factor in balancing cost efficiency with environmental sustainability. This subsection evaluates the impact of emissions caps ($\epsilon$) on the total cost structure, considering different risk-aversion levels ($\lambda$) and confidence levels ($\alpha$). The results, presented in Table 4 and illustrated in Figures LABEL:fig:OBJ_vs_E_.25, LABEL:fig:OBJ_vs_E_.5, LABEL:fig:OBJ_vs_E_.75, and LABEL:fig:OBJ_vs_E_.9, provide insights into how the emissions constraint interacts with other optimization components, shaping overall system performance.
This analysis directly addresses the second policy question on trade-offs between cost efficiency and sustainability, demonstrating how varying emissions caps impact the cost structure. Additionally, it provides insights into the first policy question by showing how decision-making under uncertainty is influenced by risk parameters when environmental constraints are introduced.
Impact of Emissions Cap on Objective Value.
Table 4 shows a clear trend: as the emissions cap increases, the objective value decreases across all configurations of $\lambda$ and $\alpha$. This behavior is a direct consequence of the constraint (9) which restricts emissions by penalizing excess emissions beyond $\epsilon$. A lower cap enforces stricter compliance, increasing the overall penalty and leading to higher costs. Conversely, a relaxed cap provides flexibility in route selection and transport decisions, reducing total costs. However, the objective function eventually stabilizes, indicating that beyond a certain point, emissions constraints cease to be the dominant cost driver.
Sensitivity to Risk-Aversion and Confidence Levels.
The model’s behavior under different risk settings highlights how emissions constraints interact with extreme cost mitigation. At higher values of $\lambda$, the optimization prioritizes risk reduction by minimizing tail-end cost scenarios, resulting in larger cost reductions when $\epsilon$ is relaxed. Similarly, higher $\alpha$ values amplify these effects, as the model becomes more sensitive to extreme cases in cost distribution. Figures LABEL:fig:OBJ_vs_E_.25LABEL:fig:OBJ_vs_E_.9 show that as risk aversion increases, the cost gap between different emissions caps widens. This reflects a trade-off: tight emissions constraints increase system resilience at the expense of higher costs.
Table 4: Objective Values for Different Emission Caps with Varying Lambda($\lambda$) and Alpha ($\alpha$) Values (scaled to $10^{3}$).
@@
Plateauing Effects and Non-Monotonic Trends.
The numerical results also exhibit plateauing behavior in objective values, particularly for higher $\epsilon$. This can be attributed to the fact that once emissions penalties no longer contribute significantly to cost, the model shifts focus to optimizing other components, such as transportation and supply costs. Furthermore, minor fluctuations in objective values at specific $\lambda$ and $\alpha$ combinations suggest potential threshold effects within the solver. These fluctuations indicate instances where the model finds alternative feasible solutions that marginally alter cost distribution but do not substantially impact overall performance.
Practical Implications for Sustainability and Cost Optimization.
The results emphasize that sustainability goals can be achieved without drastically increasing costs, provided emissions constraints are integrated strategically. At moderate $\lambda$ values, the model effectively balances emissions penalties with cost minimization, ensuring that environmental constraints do not become disproportionately restrictive. This suggests that decision-makers can set emissions limits that align with regulatory requirements without incurring excessive operational expenses.
Broader Research Implications.
This analysis provides valuable insights into the interplay between cost efficiency, emissions control, and risk preferences. It validates the model’s capacity to optimize intermodal transportation under policy-driven constraints and highlights the effectiveness of risk-adjusted decision-making in guiding sustainable logistics planning. Future research could explore dynamic emissions pricing mechanisms or the incorporation of carbon credit constraints to further enhance cost-emissions trade-offs.
@@ 5.3.6 Incorporating Intermodal Transfer Time and Cost
These constraints ensure that the total travel time, including intermodal transfers, aligns with the departure schedules of trains and prevents premature arrivals.
@@ 6 Conclusion and Future Work
This study introduces a novel two-stage stochastic optimization model for road-rail intermodal freight transportation, by integrating demand and capacity uncertainty, risk management, and sustainability objectives in a unified framework. Instead of focusing on isolated aspects of the intermodal network such as shipment allocation or carbon-efficient routing, our model uniquely combines these elements, offering a comprehensive approach to optimizing intermodal logistics under uncertainty. One of the key contributions of this study is the incorporation of real-time adaptive responses to stochastic elements such as train capacities and demand fluctuations while dynamically adjusting container allocation based on realized demand and capacity constraints. This adaptive approach ensures that decisions are aligned with operational realities, providing a level of flexibility and responsiveness that has not been previously achieved in intermodal freight optimization.
@@
Building on these unique contributions, future research could explore several promising directions. First, the model could be extended to incorporate dynamic routing decisions, allowing for real-time adjustments to traffic conditions, infrastructure disruptions, and other unforeseen events. This would further enhance the model’s adaptability and applicability in dynamic logistics environments. Second, future work could investigate the integration of advanced predictive analytics and machine learning to improve demand forecasting and capacity planning. By leveraging real-time data and predictive models, the optimization framework could be further enhanced to anticipate and respond to fluctuations in demand and capacity more effectively. Finally, the model could be expanded to include multi-objective optimization techniques that simultaneously consider additional sustainability metrics, such as energy consumption, noise pollution, and social impacts. This would provide a more holistic approach to sustainable freight transportation, addressing a broader range of environmental and social concerns.
@@
Appendix Appendix A Detailed Case Study Description
Appendix A.1 Network Description
The intermodal freight transport network for the case study used in this work consists of five major train routes connecting hubs across significant economic regions in the United States. These routes are designed to represent realistic operations in terms of distances and schedules and capacity constraints. Each hub serves as a critical transfer point for goods moving between road and rail modes, enabling efficient first-mile operations.
Appendix A.2 Example: Train A Route from Seattle to Chicago
Appendix A.2.1 Train Schedule
Train A operates on a meticulously planned schedule, departing from Seattle, WA, at 08:00 AM and arriving in Chicago, IL, at 06:00 PM on the third day. The route incorporates stops at key intermediate hubs, including Spokane, WA; Billings, MT; Fargo, ND; and Minneapolis, MN, each allowing for a six-hour dwell time. These dwell times are designed to accommodate essential freight handling operations, such as loading, unloading, and transfer activities, ensuring seamless flow throughout the network.
The detailed schedule for Train A is presented in Table A.1, capturing the precise arrival and departure times at each hub. This schedule serves as a foundation for analyzing transit times, operational efficiency, and the integration of first-mile and last-mile connectivity within the broader intermodal framework.
Appendix A.2.2 Hub-to-Hub Distances
The total distance for Train A is approximately 2,105 miles, with detailed distances between consecutive hubs provided in Table A.2. To ensure the highest level of accuracy, actual rail distances along the rail network were utilized. These distances reflect the practical realities of freight rail transport, accounting for deviations and curves inherent in rail lines compared to straight-line geodesic measurements.
By incorporating realistic rail distances, the analysis aligns closely with operational conditions, enabling precise calculations of transit times, fuel costs, and overall network efficiency. The detailed breakdown of distances between hubs in Table A.2 provides essential data for route optimization, further reinforcing the applicability of the study to real-world intermodal transport scenarios.
Appendix A.2.3 Nearby Facilities
Each intermodal hub along Train A’s route is strategically linked to nearby facilities, which act as critical supply or demand nodes for freight consolidation and distribution. These facilities, including warehouses and manufacturing plants, are selected based on their operational relevance and proximity to the hubs, ensuring seamless integration into the intermodal network.
Table A.3 provides a detailed overview of these facilities, including their names, road distances from the respective hubs, and geographical coordinates. The inclusion of these facilities not only highlights the importance of first-mile and last-mile connectivity in the intermodal system but also emphasizes the role of accurate facility-to-hub assignments in achieving cost-effective and timely freight operations. This linkage underscores the scalability and adaptability of the optimization framework in real-world applications.
Appendix A.3 Summary of All Train Routes
While Train A is discussed in detail to illustrate the methodology, the remaining four routes follow similar operational patterns, encompassing strategically selected hubs, defined schedules, and realistic distances. These routes represent a diverse set of intermodal freight connections across the United States, capturing variations in geographic coverage, transit times, and network complexity.
Table A.4 provides a comprehensive summary of the remaining routes, detailing their origins, destinations, intermediate hubs, total distances, and a high-level schedule overview. This summary serves as a concise yet informative reference, enabling practitioners and researchers to understand the broader applicability of the study. These additional routes enrich the case study by demonstrating the scalability of the proposed optimization framework to varying network configurations and operational scenarios.
Hub Departure Time Arrival Time Dwell Start Time Dwell End Time Travel Time (hours)
Buffalo, NY 12:30:00 18:57:00 19:27:00 24:57:00 6:27:00
Greencastle, PA 24:57:00 31:26:00 31:56:00 37:26:00 6:29:00
Columbus, OH 37:26:00 40:46:00 41:16:00 46:46:00 3:20:00
Cincinnati, OH 46:46:00 53:14:00 53:44:00 59:14:00 6:28:00
St. Louis, MO 59:14:00 62:24:00 62:54:00 68:24:00 3:10:00
Decatur, IL 68:24:00 75:30:00 76:00:00 81:30:00 7:06:00
Table A.1: Train A Schedule: Buffalo, NY, to Detroit, MI - Delray
Origin Hub Dest Hub Dist (m) Travel Time (hrs)
Buffalo, NY Greencastle, PA 280 6:27:00
Greencastle, PA Columbus, OH 540 6:29:00
Columbus, OH Cincinnati, OH 110 3:20:00
Cincinnati, OH St. Louis, MO 340 6:28:00
St. Louis, MO Decatur, IL 180 3:10:00
Decatur, IL Detroit, MI - Delray 310 7:06:00
Table A.2: Hub-to-Hub Distances and Transit Times for Train A
Appendix A.4 Model Parameters and Assumptions
This study employs realistic parameters for both road and rail transport to ensure accuracy and applicability. The following parameters and their ranges were used in the optimization framework:
Appendix A.4.1 Road Transport Parameters
The road segments connecting intermodal hubs to nearby facilities were modeled using an average speed of 55 miles per hour and a transportation cost of \$3.00 per mile. These values are derived from industry benchmarks for freight transport over road networks. The speed reflects variations in road types, traffic conditions, and loading/unloading delays, while the cost accounts for fuel expenses, maintenance, driver wages, and tolls. These parameters emphasize the role of facility proximity in minimizing first-mile and last-mile costs.
Appendix A.4.2 Penalty and Emission Parameters
To account for environmental and operational constraints in intermodal transport, the following parameters were included:
  • Emission Penalty ($\rho$): For exceeding emissions limits, penalties ranged from \$100 to \$500 per ton of CO2, reflecting the cost of environmental compliance and deterrence measures.
  • Unmet Demand Penalty ($\pi$): Costs for unmet demand ranged from \$1,000 to \$5,000 per container, capturing potential lost revenue and penalty fees for service failures.
  • Emission Caps ($\epsilon$): Over a 50-hour operational period, allowable emissions were capped at 50 to 150 tons of CO2, aligning with sustainability goals for the transport of 25–50 containers.
  • Carbon Emissions per Container ($E$): Emissions were estimated at 0.1–0.5 tons per container per hour, varying based on the transport mode and energy efficiency.
  • Container Preparation Costs ($O$): At origin facilities, preparation costs ranged from \$50 to \$200 per container, depending on facility efficiency and handling requirements.
Hub Facility Name Dist (m) Origin Coord (Lat, Lon) Dest Coord (Lat, Lon)
Buffalo, NY Fedex Warehouse 13 68 (43.04267661, -77.6750398) (42.8917322, -78.78303066)
Buffalo, NY Fedex Warehouse 14 16 (42.78052551, -78.84626101) (42.8917322, -78.78303066)
Greencastle, PA Fedex Warehouse 15 162 (41.2509818, -76.84433433) (39.76015595, -77.74182928)
Greencastle, PA Fedex Warehouse 16 190 (40.11206849, -74.97304753) (39.76015595, -77.74182928)
Columbus, OH Fedex Warehouse 17 23 (40.01861818, -82.9200998) (39.79411337, -82.95694498)
Columbus, OH Fedex Warehouse 18 91 (39.90979828, -84.18019515) (39.79411337, -82.95694498)
Cincinnati, OH Fedex Warehouse 19 10 (39.19316693, -84.47390985) (39.10476642, -84.53974175)
Cincinnati, OH Fedex Warehouse 20 65 (39.90979828, -84.18019515) (39.10476642, -84.53974175)
St. Louis, MO Fedex Warehouse 21 11 (38.69399342, -90.34175965) (38.68855031, -90.21324118)
St. Louis, MO Fedex Warehouse 22 29.4 (38.805481, -90.599304) (38.68855031, -90.21324118)
Decatur, IL Fedex Warehouse 23 186 (41.90215422, -87.64911463) (39.8736349, -88.91197339)
Decatur, IL Fedex Warehouse 24 79 (40.63198055, -89.45424028) (39.8736349, -88.91197339)
Detroit, MI - Delray Fedex Warehouse 25 19 (42.44998268, -83.1954107) (42.2961737, -83.12169962)
Detroit, MI - Delray Fedex Warehouse 26 28 (42.373589, -83.42067) (42.2961737, -83.12169962)
Table A.3: Nearby Facilities for Train A Route
Route Origin Destination Number of Intermediate Hubs Distance (miles)
Train B Taylor, PA New Orleans, LA 4 1,100
Train C Jacksonville, FL Ayer, MA 5 1,050
Train D Chicago 47th, IL Jacksonville, FL 4 1,200
Train E Dallas, TX Morrisville, PA 4 1,300
Table A.4: Summary of Additional Train Routes
Appendix A.4.3 Integration and Real-World Applicability
These parameters ensure a realistic representation of the trade-offs in intermodal freight transport. By incorporating penalties, emission constraints, and cost factors, the model balances operational efficiency with environmental sustainability. These considerations highlight the complexities of modern logistics, particularly in achieving carbon-neutral goals while maintaining service quality.
Appendix A.5 Reproducibility
This example provides sufficient detail to enable reproducibility and adaptation for similar studies. Researchers or practitioners can replicate the study by following these steps:
  1. 1.
    Train Schedules: Utilize detailed train schedules, including arrival, departure, and dwell times, as shown for Train A in Table A.1. These schedules ensure precise modeling of freight movements across hubs while accounting for operational constraints.
  2. 2.
    Hub-to-Hub Distances and Transit Times: Integrate hub-to-hub rail distances and corresponding transit times, as presented in Table A.2. These distances are either derived from actual rail line measurements or approximated using geodesic calculations to maintain spatial accuracy.
  3. 3.
    Facility Assignments: Assign nearby facilities to hubs based on proximity and road transport feasibility, as demonstrated in Table A.3. This ensures seamless integration within the network.
  • Train B: Taylor, PA → New Orleans, LA — NORFOLK INTERNATIONAL, VA, Charlotte, NC, Atlanta, GA, Huntsville, AL
  • Train C: Jacksonville, FL → Ayer, MA — Savannah, GA, Charleston, SC, Greensboro, NC, Pittsburgh, PA, Elizabeth Marine Terminal, NJ
  • Train D: Chicago 47th, IL → Jacksonville, FL — Louisville, KY, Georgetown, KY, Memphis, TN, Birmingham, AL
  • Train E: Dallas, TX → Morrisville, PA — New Orleans, LA, Birmingham, AL, Greer, SC, Norfolk Southern Bayview Yard, MD
While Train A serves as a detailed example in this study, summary details for the remaining four routes are also provided in the appendix (Table A.4). These summaries include key information on origins, destinations, intermediate hubs, and distances, allowing practitioners to extend the methodology to a broader intermodal network. By combining these datasets with the optimization framework described, this study ensures transparency and repeatability, laying the groundwork for future research and practical applications in intermodal freight transport optimization.

### v2

@@ 5.3.4 Impact of Train Capacity Constraints on Demand and Total Cost
Train capacity plays a critical role in intermodal transport, particularly under uncertain demand and fluctuating spot availability. To explore this, we examine a streamlined setup involving 4 freight trains, 22 intermodal hubs, and 8 demand scenarios, totaling 1013 containers (averaging 32 per train). This configuration captures key uncertainties while allowing us to evaluate how capacity expansion affects cost, demand satisfaction, and system balance.
Table 4 summarizes the results. As train capacity increases, unmet demand and total costs decrease sharply up to a point. Expanding capacity from 4 to 8 containers per station yields substantial improvements, but further increases offer only marginal gains. This identifies a capacity threshold beyond which cost savings level off, addressing our sixth policy question on the diminishing returns of capacity investment.
@@
System Performance and Trade-offs
The performance improvements stem from resolving key bottlenecks:
  • Low-capacity bottlenecks: When capacity is tight, container allocation is restricted, leading to penalties for unmet demand. Increasing capacity reduces these penalties significantly.
  • Demand volatility: With 8 demand scenarios, capacity must absorb peaks and troughs. At low levels, trains struggle to meet demand surges; higher capacities offer flexibility and reduce cost variability.
  • Diminishing returns: Beyond 8 containers, most demand is already met, and additional capacity does little to reduce costs further, as other constraints—such as supply limits and schedule alignment become more influential.
Cost vs. Unmet Demand
Figure 6 illustrates how cost and unmet demand decline with rising capacity. The steep drop between 4 and 8 containers highlights the benefit of expansion. Beyond that, improvements taper off. This plateau suggests capacity is no longer the main limiter, supply readiness, and hub synchronization are more pronounced.
The distribution of unmet demand also improves. At 4 containers, some trains face up to 18 unmet units, with a high standard deviation. By capacity 10, unmet demand is nearly eliminated and evenly distributed (maximum of 2 units, standard deviation of 0.67), reflecting a more balanced and efficient system.
@@
  • Increasing train capacity significantly improves performance but only up to a point. After 8 containers per station, gains flatten due to other operational limits.
  • At lower capacities, unmet demand varies widely across trains, indicating congestion and imbalance. Higher capacities smooth out these differences.
  • The model effectively balances cost minimization with demand satisfaction, optimizing under uncertainty.
  • Given the unpredictable nature of spot capacity (affected by external demand and network congestion), strategic planning should focus on adaptability and not just expansion.
In summary, while capacity expansion enhances flexibility and reliability, it is not a silver bullet. Beyond a certain point, improvements depend more on adaptive scheduling, coordinated resource allocation, and dynamic response to real-time constraints.
@@ 5.3.5 Sensitivity of Cost Metrics to Risk Preferences and Scenario Variability
A central challenge in stochastic intermodal planning is balancing cost efficiency with robustness across uncertain scenarios. While previous sections explored how $\lambda$ and $\alpha$ shape total cost and CVaR, this subsection focuses on the model’s sensitivity to cost variability and stability. Specifically, it examines the divergence between the objective value (OBJ) and expected total cost $E(TC)$, the distribution of costs across scenarios, and how these factors shift with increasing risk aversion. Figures 7 and 8 visualize these dynamics.
This analysis informs Policy Questions 1, 3, and 7 by illustrating the implications of risk preferences on cost predictability and revealing the diminishing returns of extreme risk aversion.
@@
Figure 7 shows that as $\lambda$ increases, particularly at higher $\alpha$, the gap between OBJ and $E(TC)$ widens. At low $\lambda$, both metrics remain closely aligned, suggesting that cost minimization dominates the objective function. However, as $\lambda$ grows, CVaR becomes more influential, indicating a shift in the model’s focus toward managing worst-case outcomes. This divergence captures the cost-resilience trade-off inherent in conservative planning.
Scenario-level variability, illustrated in Figure 8, further confirms this trend. At low $\lambda$, the model achieves lower average costs but with significant fluctuation across scenarios. As risk aversion increases, cost volatility declines, producing more stable financial outcomes. Yet, this stabilization levels off around $\lambda\approx 0.8$, beyond which additional risk mitigation yields marginal improvements in variability while significantly raising total cost. This plateau suggests that moderate values of $\lambda$ provide the most efficient balance between stability and cost.
The CVaR sensitivity analysis reinforces these observations. Increasing $\lambda$ reduces CVaR, but the rate of reduction diminishes as $\lambda$ approaches 0.5. The impact of risk aversion is more pronounced at higher $\alpha$ (e.g., $\alpha=0.90$), where tail-end scenarios carry greater weight. These findings indicate that moderate risk preferences typically $\lambda$ between 0.4 and 0.6 offer strong protection against extreme costs without incurring excessive expense.
@@
Figures LABEL:fig:ETC_surface and LABEL:fig:ETC_contour map $E(TC)$ across the $(\lambda,\alpha)$ space, highlighting how predictable cost structures emerge at intermediate risk settings. At low $\lambda$, the model aggressively allocates costs for efficiency, but at the expense of stability. At high $\lambda$, cost structures flatten but total costs surge, making the trade-offs more pronounced.
From a practical standpoint, results emphasize the need to tailor risk settings to organizational priorities. Low-risk ($\lambda\leq 0.3$) may suit stable environments but carry vulnerability to high-impact scenarios. Balanced settings near $\lambda=0.5$ offer predictable costs with modest premiums, while highly conservative plans ($\lambda\geq 0.8$) deliver resilience at major cost. This flexibility enables stakeholders to calibrate the model according to their risk tolerance and operational needs.
These findings underscore the importance of aligning model risk settings with strategic objectives. They also highlight potential for exploring how risk preferences interact with emissions penalties, capacity constraints, or real-world disruptions like terminal congestion and supply bottlenecks.
@@ 5.3.6 Analyzing the Influence of Emissions Cap on Model Performance
Emissions constraints are a critical lever in aligning cost efficiency with sustainability objectives in intermodal transport. This subsection evaluates the impact of emissions caps ($\epsilon$) on total cost under varying risk aversion levels ($\lambda$) and confidence levels ($\alpha$). The results, presented in Table 5 and visualized in Figures LABEL:fig:OBJ_vs_E_.25 through LABEL:fig:OBJ_vs_E_.9, reveal how the emissions constraint influences system-wide performance and interacts with the model’s broader cost structure.
As shown in Table 5, increasing the $\epsilon$ consistently reduces the objective across combinations of $\lambda$ and $\alpha$. This trend stems from the emissions constraint defined in Equation (10), which imposes penalties for exceeding $\epsilon$. Stricter limits (lower $\epsilon$) drive up costs by restricting routing flexibility and prioritizing lower-emission options, while relaxed limits ease those penalties and allow more cost-effective routing. However, the cost benefit of relaxation eventually plateaus, indicating that emissions constraints stop being the primary cost driver once penalties fall below a certain threshold.
Table 5: Objective Values for Different Emission Caps with Varying Lambda($\lambda$) and Alpha ($\alpha$) Values (costs in \$ $\times$ 103)
@@
The interaction between emissions constraints and risk preferences is also notable. At higher $\lambda$ values, the model places greater emphasis on mitigating extreme costs, making it sensitive to changes in $\epsilon$. This effect intensifies with higher $\alpha$, which shifts focus toward tail-end scenarios. As illustrated in Figures LABEL:fig:OBJ_vs_E_.25LABEL:fig:OBJ_vs_E_.9, the cost gap between different emissions caps widens with increasing risk aversion. This demonstrates a clear trade-off: tighter emissions limits can enhance system resilience but at the cost of elevated expenditures, particularly in more conservative planning regimes.
Interestingly, the objective values begin to plateau at higher emissions thresholds, and in some cases, exhibit minor non-monotonic fluctuations. These irregularities shows solver behavior near optimality, where multiple feasible solutions deliver similar outcomes but slightly redistribute cost components such as supply or transport. Once emissions penalties become negligible, the model’s focus naturally shifts toward optimizing other drivers.
These findings confirm that substantial emissions reductions can be achieved with minimal cost trade-offs, particularly when emissions constraints are calibrated effectively. At moderate $\lambda$ and $\alpha$ values, the model jointly balances risk, cost, and sustainability objectives, supporting operational decisions that align with environmental policy. Notably, the results indicate a threshold effect: beyond an emissions cap of approximately 200–250 tons, further relaxation yields little marginal cost benefit. This suggests that carbon caps or pricing schemes targeting this range can serve as effective policy levers, encouraging mode shifts without imposing excessive cost burdens. These insights offer practical value for regulators or enterprises evaluating carbon tax thresholds or emissions allowances, and they support future extensions involving explicit carbon pricing, credit trading mechanisms, or adaptive policy rules.
@@ 5.3.7 Incorporating Intermodal Transfer Time and Cost
These constraints ensure that travel time, including intermodal transfers, aligns with trains schedules.
@@ 6 Conclusion and Future Work
This study introduces a novel two-stage stochastic optimization model for road-rail intermodal freight transportation, by integrating demand and capacity uncertainty, risk management, and sustainability objectives in a unified framework. Our model uniquely combines aspects of the intermodal network such as shipment allocation and carbon-efficient routing, offering a comprehensive approach to optimizing intermodal logistics under uncertainty. One of the key contributions of this study is the incorporation of real-time adaptive responses to stochastic elements such as train capacities and demand fluctuations while dynamically adjusting container allocation based on realized demand and capacity constraints. This adaptive approach ensures that decisions are aligned with operational realities, providing a level of flexibility and responsiveness that has not been previously achieved in intermodal freight optimization.
@@
Building on these contributions, future work could extend the model to incorporate dynamic routing and variable travel time. In this study, we assumed fixed road transport times for tractability and consistency with short-haul, repetitive intermodal routes often assumed to be fixed (Basallo-Triana et al.,, 2021). The intermodal network structure and routing paths were also predetermined. While these assumptions are suitable for strategic planning, they limit responsiveness to real-time disruptions such as congestion or weather events. Acknowledging these limitations, future work could explore real-time optimization or simulation-based methods to enhance the model’s adaptability to dynamic conditions. Additionally, since we used a commercial solver, exploring heuristic or decomposition-based approaches may improve scalability for large real-world instances. Second, future work could investigate integration of predictive analytics and machine learning to improve demand forecasting and capacity planning. By leveraging real-time data and predictive models, the optimization framework could be further enhanced to anticipate and respond to fluctuations in demand and capacity more effectively. Finally, the model could be expanded to include multi-objective optimization techniques that simultaneously consider additional sustainability metrics, such as energy consumption. This would provide a more holistic approach to sustainable freight transport, addressing a wider range of environmental concerns.
@@
Acknowledgment
This work was partially supported by the U.S. Department of Energy’s Advanced Research Projects Agency-Energy (ARPA-E) under the project (#DE-AR0001780) titled “A Cognitive Freight Transportation Digital Twin for Resiliency and Emission Control Through Optimizing Intermodal Logistics".